# Learning phases with Quantum Monte Carlo simulation cell

### v1

@@ V Conclusions
In conclusion, based on Stochastic Series Expansion QMC we have proposed a new input data type “spin-opstring” for machine learning applications. The spin-opstring is a compact representation of the QMC simulation cell that retains all quantum information of the system while being memory-efficient. It consists of the initial state followed by an operator string that describes the state’s evolution through imaginary time, ultimately bringing it back to its original configuration. In order to demonstrate the applicability of spin-opstring as ML input, we have applied supervised machine learning on two different models : the two-dimensional $t_{2}-V_{1}$ model and the one-dimensional Su-Schrieffer-Heeger model. We found that in both models the input data performs very well by capturing the phase transitions accurately. This highlights the versatility of our input data in identifying conventional phase transitions, both continuous and discontinuous, as well as topological phase transitions. Moreover, we employed the spin-opstring to perform a regression task predicting the superfluid density in the $2D$ $t_{2}-V_{1}$ model. These results provides additional evidence for the suitability of our proposed input data in machine learning applications. It would be interesting to see how the spin-opstring data perform in the domain of unsupervised ML algorithms.

### v2

@@ VIII Conclusions
In conclusion, we have proposed to use the “spin-opstring”, extracted from Stochastic Series Expansion QMC as a input data type for machine learning applications. The spin-opstring is a compact representation of the QMC simulation cell that retains all quantum information of the system while being memory-efficient. It consists of the initial state of the simulation cell followed by an operator string that describes the state’s evolution through imaginary time, ultimately bringing it back to its original configuration. In order to demonstrate the applicability of spin-opstring as ML input, we have applied supervised machine learning on two different models : the two-dimensional $t_{2}-V_{1}$ model and the one-dimensional Su-Schrieffer-Heeger (SSH) model. We found that in both models the input data performs very well by capturing the phase transitions accurately. This highlights the versatility of our input data in identifying conventional phase transitions, both continuous and discontinuous, as well as topological phase transitions. Moreover, we employed the spin-opstring to perform a regression task predicting the superfluid density in the $2D$ $t_{2}-V_{1}$ model. We have further compared the performance of the spin-opstring with the full simulation cell data. Our results show that as system size increases, the combined memory and computational costs make the full simulation cell impractical.
To explore the broader applicability of the spin-opstring, we conducted transfer learning experiments across different physical models. First, a model trained on the $t_{2}-V_{1}$ system was successfully applied to predict phase transitions in the $t_{1}-V_{1}$ model which features a similar superfluid–checkerboard solid transition but originates from nearest-neighbor (NN) hopping rather than next-nearest-neighbor (NNN) hopping. Despite this difference in origin, the model can very well apply the learned knowledge from the $t_{2}-V_{1}$ model to identify the phase boundary in the new system. In a second case, a ML model trained on the non-interacting SSH model accurately predicted phase transitions in the interacting SSH model with nearest-neighbor repulsion, up to some interaction strength. These results illustrate the strong potential of spin-opstring data in facilitating knowledge transfer between related models. We also evaluated generalization across system sizes, by training a ML network on the SSH model in a smaller lattice and testing it on a larger system. The model correctly identified the topological-to-trivial phase transition, confirming that spin-opstring data supports generalization to larger systems.
Overall, our results establish that the combined spin-opstring can successfully be applied across a wide range of ML tasks. For completeness we also addressed the question whether either the spin-state or the operator string could serve as stand-alone inputs. We considered examples to demonstrate that the unified spin-opstring representation leads to a more reliable model performance than using either component individually.
Finally, we employed two leading interpretability techniques, Layer-wise Relevance Propagation (LRP) and SHapley Additive exPlanations (SHAP), to gain insight into the model’s decision-making processes. We devised a strategy to make the high-dimensional outputs of LRP and SHAP interpretable by linking the feature scores to specific physical operators. Our analysis revealed that while LRP disregards identity operators (as they are encoded as zeros in the spin-opstring), SHAP is able to recognize their contextual importance. This highlights the advantage of SHAP in scenarios where every component of the input, including identity operators, contribute to the model’s predictions. Across classification, regression, and transfer learning tasks, our interpretability study consistently showed that the models learned physically meaningful features from the input data. This supports the conclusion that the spin-opstring is not only effective but also interpretable for ML-based analysis of quantum systems.
In summary, our results establish the spin-opstring as a broadly applicable, generalizable, and interpretable input format for machine learning in quantum many-body physics. It would be interesting to see how the spin-opstring data perform in the domain of unsupervised ML algorithms. Another interesting direction of our future research involves the use of generative ML models with the spin-opstring data representation for the generation of SSE configurations.
@@ Appendix A Effect of Batch Normalization
BN layers are commonly applied to ML networks to mitigate “internal covariate shift”, a phenomenon where the distribution of inputs to a layer in a neural network changes during training due to the continuous updates of parameters in preceding layers. In our spin-opstring data derived from SSE QMC simulations, the operator string segments may exhibit considerable lengths (e.g., up to $5\times 10^{4}$ elements for $12\times 12$ lattices in the $2D$ $t_{2}-V_{1}$ model) and substantial fluctuations for each element, $S(p)$. As argued in Ref. [43], the absence of BN in such scenarios can lead the ML model to overfit and be susceptible to the initial conditions of the training process. This is consistent with our observation of a wide range of predictions for the phase transition point in Fig. 23 top.